OpenStudy (anonymous):

What is the value of x? 3(x-2)+5x=18

OpenStudy (anonymous):

@iGreen @AlexandervonHumboldt2

OpenStudy (igreen):

Distribute 3 into the parenthesis first.. \(3(x - 2) \rightarrow (3 \cdot x) + (3 \cdot -2)\)

OpenStudy (anonymous):

so the answer is -6?

OpenStudy (igreen):

No, we're not done yet..do you know what \(3 \cdot x\) is?

OpenStudy (anonymous):

no

OpenStudy (igreen):

\(3 \cdot x \rightarrow 3x\) So we get \(3x - 6\). Plug that back in: \(3x - 6 + 5x = 18\) Now can you add 3x + 5x?

OpenStudy (anonymous):

so 8x

OpenStudy (anonymous):

8?

OpenStudy (anonymous):

3-6=-3

OpenStudy (anonymous):

so it would be -3x + 5x=18

OpenStudy (anonymous):

idk i am so confused

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@AnswerMyQuestions

OpenStudy (anonymous):

Where exactly are you confused?

OpenStudy (anonymous):

srry on everything

OpenStudy (anonymous):

plz explain it to me

OpenStudy (anonymous):

 So I think this is the part that can get most confusing, and hopefully this makes sense, now you want to move all the terms with x to one side and all the integers to the other (collecting like terms). In these sorts of questions, when we do algebra, we want to isolate the variable (the letter), here it's x. So what you do to one side you must also do to the other, and to move things around you do the opposite operation of what already exists. So for example: 4x+2=2 We want to isolate "x" get it on its own. Our first step is to move the 2 on the left side to the right, to do this we say alright, so how can that be done, we use the opposite operation, subtract 2 on both sides. So now you'll have 4x+2-2=2-2 which is 4x+0=0 4x= 0 now we say ok how do we move the 4 on the right side since we want x alone. You see the 4 is "stuck" onto the x meaning it's being multiplied, so once again opposite operation would be divide. Divide both sides by 4 now. We get 4x/4 = 0/4 Which is x = 0, hope that makes some sense!